Find the LCM of 6x^8 and 12x^5
Answer by stanbon(57222)

Find the LCM of 6x^8 and 12x^5
-----
The LCM is the product of all the factors of the component terms
in their highest power.
6x^8 = 2*3*x^8
12x^5 = 2^2*3*x^5
----------
So the LCM must have a 2^2; a 3; and x^8
So the LCM is 2^2*3*x^8= 12x^8
=============
cheers,
Stan H.